OpenStudy (anonymous):

what is (-2/sqrt(x))^(6)

OpenStudy (anonymous):



OpenStudy (anonymous):

then simplify numerator (-2)^6 = --------------? & denominator ( X^1/2)^6 = x^3

OpenStudy (anonymous):

Do you get it Kayla ?

OpenStudy (anonymous):

Thanks! Now I have to multiply it with 210 . x^4

OpenStudy (anonymous):

It was a binomial expansion problem

OpenStudy (anonymous):

:D